### 16 Feb 2024 v1

# Tailoring Polarization in WSe Quantum Emitters through Deterministic Strain Engineering

Abstract
A quantum emitter in a transition metal dichalcogenide has recently emerged as a promising platform for generating single photons for optical quantum information processing. In this work, we present an approach for deterministically controlling the polarization of fabricated quantum emitters in a tungsten diselenide (WSe) monolayer. We employ unconventional nanopillar geometries with high apex-to-base ratios and sharp vertices to induce a controlled directional strain in the monolayer, and we report on fabricated WSe emitters producing single photons with a high degree of polarization and high purity (). Our work paves the way for the deterministic integration of TMD-based quantum emitters in quantum photonic architecture.
I Introduction
Within photonic quantum information technologies, a key component is the source of single photons Gregersen et al. (2017) used to encode the quantum bits. The ideal single-photon source (SPS) should feature deterministic emission of single indistinguishable photons with near-unity collection efficiency and high purity. The spontaneous parametric down-conversion process is a straightforward method Kwiat et al. (1995) for producing highly indistinguishable photons and has been the workhorse of the community for several decades; however, its probabilistic nature results in a significant trade-off between photon purity and efficiency. As an alternative, the two-level system in a semiconductor host material Aharonovich et al. (2016) has recently emerged as an attractive platform for the deterministic generation of single photons using the spontaneous emission process. By placing a semiconductor quantum dot (QD) in an optical cavity Heindel et al. (2023) and by exploiting cavity quantum electrodynamics (cQED) to control the light emission, emission of pure highly indistinguishable photons with an efficiency of 0.6 has been demonstrated Tomm et al. (2021); Wang et al. (2019). However, the fabrication of high-quality semiconductor QDs requires expensive unreliable molecular beam epitaxy growth methods representing a major drawback in the development of QD-based sources.
As an alternative, quantum emitters in transition metal dichalcogenides (TMDs) are attractive due to their availability, their ease of integration into nanophotonic structures, and the versatility in engineering their photonic characteristics Montblanch et al. (2023). While the direct bandgap structure Le et al. (2015); Montblanch et al. (2023) of the pristine monolayer form allows TMDs to act as an efficient classical light source, the emission of pure single photons requires additional engineering. The microscopic origin of single-photon emission in TMDs is believed to be localized excitonic states Aharonovich et al. (2016); Gao et al. (2023); Iff et al. (2019); Linhart et al. (2019); Parto et al. (2021); Branny et al. (2017) appearing either due to defects in the crystal lattice or due to the application of strain. Thus the implementation of single-photon emitters (SPEs) into TMDs has been pursued by using deterministic defect fabrication Parto et al. (2021); Xu et al. (2023); Wang et al. (2022) as well as by introducing localized strain in TMD monolayers Kumar et al. (2015); Kern et al. (2016). While stress can be applied directly to the flat TMD lattice Iff et al. (2019); So et al. (2021); Xu et al. (2023); Yu et al. (2021), the strain is typically introduced by placing the TMD layer on arrays of nanopillars Parto et al. (2021); Iff et al. (2021); Branny et al. (2017); Palacios-Berraquero et al. (2017); Kim et al. (2019); Zhao et al. (2021); Luo et al. (2018); Azzam et al. (2023); Kern et al. (2016) leading to strain concentration at the contact point.
Similar to QDs, efficient single-photon emission can be obtained from an SPE in a TMD by placing it inside an optical cavity and exploiting cQED. A WSe emitter was recently placed inside an open cavity geometry Drawer et al. (2023) leading to the demonstration of an efficiency of 0.65 as well as 0.02 indistinguishability. The cQED effect relies on a matching of polarizations of the dipole of the emitter and the electric field of the optical cavity mode. For the strain-induced SPE, a deformation strain potential, a nanowrinkle, serving as a macroscopic host for the SPE, is produced by the contact of the WSe monolayer with the nanopillar. The polarization of the SPE dipole is then aligned So et al. (2021); Azzam et al. (2023); Wang et al. (2021) with the orientation of the wrinkle. Control of the lateral dipole orientation is not needed in a rotationally symmetric geometry Drawer et al. (2023) with two degenerate orthogonally polarized optical cavity modes, where the dipole can couple into either mode. However, the general case requires control of the SPE dipole polarization, and thus the wrinkle orientation, to efficiently couple light into, for example, the polarized TE mode of a planar ridge waveguide as needed for on-chip integrated quantum photonics Wang et al. (2020).
In this work, we propose a simple and efficient method for fabricating orientation-controlled wrinkles using strain-engineered monolayer WSe, and we demonstrate light emission from SPEs with high purity () and a high degree of polarization (99 4%). Our method is based on nanostructures tailored to deterministically induce strain in a specific direction. Moreover, we artificially introduce lattice defects with electron beam irradiation to further enhance the emission process. This work paves the way for the controlled integration of TMD-based quantum emitters into photonic systems.
II Results and Discussions
II.1 Engineering the WSe quantum emitters via cylindrical nanopillars
The material used in this study, WSe, has an emission wavelength range (750-800 nm) compatible with low-loss free-space channels and with the wavelength of Rb-87-D2 for integration with quantum memories Drawer et al. (2023); Simon et al. (2010).
We initially fabricated single-photon emitters on conventional Parto et al. (2021); Palacios-Berraquero et al. (2017) rotationally symmetric nanopillars. Pristine monolayer flakes of WSe are transferred on top of an array of cylindrical nanopillars using the dry transfer technique Castellanos-Gomez et al. (2014). The pillar dimensions were set to 150 nm in height and 0.5 µm in diameter. To adapt to the shape of the pillar, the flakes undergo mechanical strain, deform, and create nanowrinkles. The detailed fabrication process is discussed in the Methods section as well as in Figure S1 and Supporting Note 1. Bright field (BF) and photoluminescence (PL, T = 4 K) images of the fabricated sample with cylindrical pillars are shown in Figures 1a,b, respectively. In Figure 1a, a white outline highlights the monolayer, while the pillars are identified through their contrast difference, appearing as dark spots. As can be seen in Figure 1b, a brighter PL emission signal surrounding the pillar was observed, in comparison with the planar region of WSe. This PL increase could be attributed to the efficient migration of excitons toward lower-energy states in the strained areas. Interestingly, the brighter PL emission signal seems to be located around the pillar. This suggests that a stronger strain is localized around the pillar, not on top, which is in good agreement with the literature Branny et al. (2017). However, it is unclear from the PL image how the nanopillar induces a local strain in the monolayer.
To provide microscopic details of the deformation around the pillar, we performed the scanning electron microscope (SEM) imaging of the fabricated sample, presented in Figure 1c. The obtained SEM image reveals the formation of four straight nanowrinkles around the pillar.
It is generally accepted that these nano-wrinkles host quantum emitters due to local strain, and their orientation is what dictates the polarization profile of the emitted single photons from the resulting emitters So et al. (2021); Azzam et al. (2023); Wang et al. (2021).
Due to the rotational symmetry of the nanopillar, which applies strain uniformly, the final number and orientation of these nanowrinkles are random, rendering the results of each sample irreproducible. Moreover, their close proximity and difficult spatial localization of nanowrinkles around the nanopillar could potentially affect further optical characterizations, e.g., purity.
To investigate the impact of closely formed nanowrinkles in the presence of a single pillar on the optical properties of the sample, we measured PL spectra (T = 4 K) from five different spots around the pillar, highlighted with white crosses (S1-S5) and present them in Figure 1d. A broad emission signal is observed from the planar monolayer region (S1) due to the delocalized excitons, whereas the collected spectra taken from different spots around the nanowrinkles (S2-S5) exhibit new discrete and narrow emission lines with higher intensities ranging from 730 - 810 nm. These narrow emission lines are attributed to the new localized excitons hosted in the confined systems produced by the strain and defects in the mono- and a few layers WSe Montblanch et al. (2023); Dang et al. (2020); Xu et al. (2023); Linhart et al. (2019); Parto et al. (2021).
The broad emission, similar to the planar monolayer, was also observed from the nanowrinkles (S2-S5) in the spectral range between 720-770 nm due to the collection spot area (larger than the emitter region), which leads to partial collection from the planar region. This makes isolating individual emission lines associated with potential single-photon emission from the broad multiphoton emission background relatively challenging, even with narrow optical bandpass filters. However, since these narrow lines appear relatively intense, it still might be possible to perform further characterization. Moreover, due to the spatial proximity of the nanowrinkles, the PL spectra exhibit overlapping emission lines, as we capture light emitted from multiple emitters simultaneously. Furthermore, these features might share their in-plane polarization axis if the nanowrinkles hosting them have a parallel orientation (S2-S4, S3-S5), making their spectral isolation practically impossible. Nevertheless, we found several intense individual emission lines in longer wavelengths above 780 nm, which could be easily isolated using a long-pass optical filter. One crucial characteristic of quantum emitters, highly relevant to numerous quantum photonic applications, is the degree of purity, indicating whether they emit only one photon at a time. This single-photon purity is often assessed through the value. To assess the purity of these new emission lines in Figure 1d, we performed the second-order autocorrelation () measurement from a representative emission line ( nm) under continuous-wave (CW) excitation at 455 nm via a Hanbury–Brown–Twiss (HBT) setup. In this setup, the emitted photons are coupled to a 50:50 fiber beam splitter, whose output signals are detected by superconducting nanowire single-photon detectors (SNSPD). The emission line at 807 nm was spectrally isolated using a 750 nm long-pass filter. Figure 1e shows the obtained correlation histogram, exhibiting a clear signature of photon antibunching. The extracted value of is below the threshold, indicating single-photon emission behavior. Despite the apparent successful fabrication of SPEs with cylindrical pillars, the challenges posed by the irreproducibility of results and the lack of deterministic control of polarization in TMD-based emitters underscore the need for an effective strain-engineering approach to potentially overcome these limitations. In the following section, we discuss a unique method of employing novel nanostructure designs to induce directional strain in the WSe monolayer in a deterministic manner for controlling the in-plane polarization and spatial localization of the WSe emitters.
II.2 Novel nanopillar designs
To address the aforementioned limitations, we propose unique nanostructure designs with tailored geometries for deterministic strain engineering to fabricate highly-polarized SPEs from monolayer WSe. The proposed pillar designs consist of polygonal structures with long vertices, sharp edges, and a relatively high apex-to-base ratio. Their shape allows us to induce the strains on the tips of the vertices and facilitate the creation of spatially isolated nanowrinkles with controlled orientation, which follows the directionality of the vertices (see Figure 2a). Controlling the direction of these one-dimensional nanowrinkles, which may host emitters, gives rise to an effective manipulation of the emitted linearly-polarized single photons So et al. (2021); Azzam et al. (2023); Wang et al. (2021). Here, we explore three different pillar geometries, namely a three-pointed star (TS), a five-pointed star (FS), and a bowtie (BT). The structures are patterned out of negative electron beam resist (HSQ), which is cured into SiO upon e-beam exposure. The thickness of the nanostructures ( nm) was identified as optimal for producing nanowrinkles during the transfer process. Further details of the e-beam lithography and the remaining fabrication process are described in the Methods section.
First, we demonstrate WSe single-photon emitters with a TS pillar. Figure 2a shows a 3D schematic representation of the TS nanostructure with a WSe monolayer flake, illustrating the formation of nanowrinkles that follow the directionality of each vertex. The surface morphology of a sample hosting a WSe monolayer flake is shown in the atomic force microscope (AFM) image (see Figure 2b). The color bar accompanying the AFM image highlights the height of the TS structure while emphasizing the height difference between the fabricated nanowrinkles and the planar monolayer. We examine the morphological profiles of the fabricated nanowrinkles on this exemplary structure in Figure S3 of the SI. As shown in Figure 2b, the monolayer forms nanowrinkles mostly following the directionality of the vertices. Moreover, the nanowrinkles are separated by more than 2 µm, allowing us to address each emitter individually.
To investigate how these nanowrinkles modify the optical properties of monolayer WSe, we performed PL imaging under a 470 nm light-emitting diode (LED) excitation in a closed-cycle optical cryostat equipped with a microscope objective (60, NA = 0.82) operating at T=4 K.
The obtained PL image (T=4 K) taken from the TS sample is presented in Figure 2c, where increased PL emission is observed along the nanowrinkles. This points to the existence of single-photon emitters due to the formed directional strains in the sample.
The emission in the area of the TS structure, highlighted by a semi-transparent star, is comparable to that of the planar monolayer, confirming that it is solely produced by the nanowrinkles. More notably, the increased PL signal is not only focused on the tips of the vertices but can also be seen across the length of the nanowrinkles.
To gain further insights into the effect of the formed nanowrinkles on the polarization of emitted photons from the spatially separated nanowrinkles in the TS sample, we performed the polarization-resolved µPL measurement from the nanowrinkles using a confocal microscope setup.
The collected PL spectra from the individual nanowrinkles (P1-P3) under parallel and orthogonal polarizations with respect to the directionality of each nanowrinkle are presented in Figures 2d-f. They exhibit highly polarized emission, as the PL intensity of the orthogonal polarization is orders of magnitude lower than the parallel one. This behavior implies that the polarization orientation follows the directionality of the nanowrinkles.
To gain a better understanding, we analyzed the polarization-resolved µPL spectra taken from the individual nanowrinkles. The extracted PL intensity as a function of the in-plane polarization exhibits highly linear polarization behavior, as shown in Figure 2g-i. To quantify the degree of linear polarization (DOLP) of the emitted photons, we use a sinusoidal fitting function and extract the DOLP using the maximum and minimum PL intensities as , with the calculated values of 99 ± 4%, 82 ± 14%, and 93 ± 3%, respectively. The strong linear polarization is attributed to the nanowrinkles, which give rise to a localized one-dimensional strain potential, consequently forming strongly aligned dipoles. Tripathi et al. (2018); Drawer et al. (2023) In addition, the fitting reveals that the in-plane polarization profile of each emitter follows the directionality of their respective nanowrinkles.
Based on these results, it is evident that the proposed design facilitates the successful isolation of the emitters both spatially and in terms of polarization.
More importantly, our method indeed enables reproducible and effective control of the in-plane polarization of the emitted photons by deterministically inducing directional strain in the nanowrinkle with sharp vertices.
Figure 3, as in the TS case, presents a similar investigation of the single photon emission property of WSe on the FS structure. The AFM image in Figure 3 showcases the surface morphology of the FS sample and the resulting nanowrinkles. The false-shading representation is shown in the inset for better contrast. We notice that one vertex out of the five vertices does not have a visible nanowrinkle forming at its end. This is dismissed as a normal occurrence due to the intricacies of the transfer process. Moreover, we observe that in the other four vertices (P1-P4) the nanowrinkles are significantly shorter than that of the TS, which is attributed to the increased number of vertices in the FS that may obstruct the formation of longer nanowrinkles. Nevertheless, the nanowrinkles are well separated and aligned with the vertices.
To further analyze the optical properties of the monolayer WSe on the FS structure, we obtain the corresponding PL image ( K) under a 470 nm LED excitation, as shown in Figure 3b. The PL signal significantly increases around the tips where the nanowrinkles are localized (P1-P4), with the tip missing a nanowrinkle showing emission comparable to the pristine monolayer. While we understand that the presence of SPEs relies on the one-dimensional potential induced via strain engineering, it becomes evident that the formation of a nanowrinkle is a prerequisite.
To understand the effect of the nanowrinkles of the FS structure on the in-plane polarization of the single photons and compare it to that of the TS, we obtain polarization-resolved µPL spectra from the four highlighted spots. We extract the PL intensities and present them as polar plots in Figures 3c-f. Similarly to Figure 2, the sinusoidal fitting reveals that the emitted photons from each SPE demonstrate high DOLP, with the extracted values at 86 , 81 , 76 , and 85 , respectively. More importantly, the orientation of the in-plane polarization follows the directionality of the vertices, verifying our ability to control the polarization of the SPEs.
While these structures demonstrate promising control over both the SPEs’ localization and polarization, it is worth emphasizing that the control of wrinkle directionality is not flawless. As shown in Figures 2 and 3, possible misalignments between the vertex and the nanowrinkle are possible. This behavior is attributed to the hexagonal lattice structure of the material and how its intrinsic “armchair” and “zigzag” directions Le and Rahman (2013) might affect the final directionality of the nanowrinkles. Statistical analysis of the deviation in alignment between the directionality of the vertices and the nanowrinkles of the TS and FS structures is presented in Supporting information Note S4 and Figure S4. Despite this analysis indicating that there are no overbearing misalignment issues, an upgraded design is presented in Figure 4. The BT design facilitates the formation of the nanowrinkle between the two triangles. This way, any potential misalignment between the vertices and the nanowrinkle would be minimized while isolating the SPEs from any naturally forming deformation in the surrounding area. Figure 4a shows the AFM image of an exemplary BT structure hosting the monolayer WSe flake. A clear nanowrinkle connecting the two triangles is visible.
Figure 4b presents the corresponding PL image of the structure, showing the increased photon emission between the nanowrinkle (Q1). Furthermore, increased photon emission is also observed in the left corner of the top triangle, which is attributed to a nanowrinkle formed along the edge of the structure. Evidently, the separation between the two regions is sufficient to characterize them individually.
Additionally, Figure S5 presents four different SEM images of exemplary bowtie structures, all hosting a nanowrinkle connecting the two triangles, with a bar graph illustrating the general trend of directionality. Even though a statistical analysis is not possible in such a small sample, the trend is obvious, as the ’misaligned’ nanowrinkles do not deviate by more than three degrees.
After examining all suggested structural designs, we have demonstrated significant control over the polarization of the produced emitters and their spatial isolation. These findings provide crucial insights for the future integration of these quantum emitters into a specific mode of desired photonic structures, e.g., the TE mode of a planar ridge waveguide. Understanding key aspects of deterministic polarization control of quantum emitters and fabrication limitations paves the way for their effective utilization as functional single-photon sources in quantum photonics.
II.3 Characterization of single-photon emission
In order to assess the quality of the fabricated SPEs based on the novel structures presented in this work, we proceed to the optical characterization of an exemplary WSe nanowrinkle quantum emitter. To perform µPL measurements, we employ a 532 nm femtosecond pulsed laser (80 MHz) to optically excite the emitter. Figure 5a shows a broad range (600-900 nm) µPL spectrum of an exemplary WSe nanowrinkle emitter with a central wavelength of 806.1 nm. A 750 nm long-pass filter was used to spectrally isolate the single-photon emission line from the delocalized exciton emission and excitation laser. The inset of Figure 5a shows a high-resolution PL spectrum of the same emitter around its central wavelength, which exhibits a prominent zero-phonon line (ZPL) with a lower-energy phonon side band (PSB). The spectrum is fitted using a sum of a Lorentzian, representing the ZPL contribution (dashed red line), and a Gaussian, which captures the information related to the PSB contribution (dashed green line). We extract the linewidths of the two fitted curves, giving values of nm and nm, respectively, with the central wavelength of the PSB being red-detuned by 0.6 nm from the ZPL, resulting in an asymmetric PL emission spectrum. The increased contribution of the PSB to the total emission, evidenced by the ratio between the intensities ZPL : PSB = 49.2 % : 50.8 %, suggests a strong exciton-phonon coupling for WSe monolayer SPEs. Mitryakhin et al. (2023) Even though the linewidth of the emission line is not Fourier-limited, being able to attain a narrow ZPL and resolve the PSB is attributed to the fabrication method that results in improved spatial isolation and a high linearly polarized emission. In addition, the spectral diffusion of the emission lines can also be affected under different excitation schemes. The excitation of the monolayer by a CW LED introduces thermal fluctuation and severe charge noise around the emitter due to photo-excited free carriers, leading to additional broadening of the emission line. To visualize the effect of additional broadening under strong CW LED excitation, we measured PL spectra under CW 470 nm LED and 532 nm pulsed laser excitation from the same emitter (see Supporting Note 5 and Figure S6). Indeed, the obtained PL spectrum under CW LED excitation exhibits a much broader and red-shifted emission line, in contrast to the pulsed excitation. To obtain the decay time of the exemplary emitter, we perform a time-resolved PL measurement. The result is presented in Figure 5b. Two different decay dynamics are clearly visible from the plot and are fitted with two separate exponential functions. The faster one exhibits a radiative lifetime at ns while the slow one is at ns. The slow decay (purple fit in 5b) is the one that dominates the time-resolved response, making the greatest contribution to the overall curve, and can be credited to the single-photon emission. Moreover, the lifetime ranges reported from previous works with regard to localized SPEs on WSe monolayer on dielectric substrates Kumar et al. (2015); Tripathi et al. (2018); So et al. (2021); Iff et al. (2021); Branny et al. (2017); He et al. (2015); Srivastava et al. (2015) agree with the reported results. The fast decay may be attributed to the instrument response function (IRF), while the abnormal increase in counts before the second decay begins may be caused by several factors. Trap states or excitation transfer processes could lead to delayed decay and an initial rise in lifetime. Concurrently, non-radiative pathways, potentially involving interactions with defects, may compete with radiative decay, contributing to this rise. Additionally, exciton dynamics, including formation and migration processes, may also influence the lifetime dynamics.
To study the stability of the emission signal from the fabricated WSe nanowrinkle, we obtain the high-resolution time trace of the PL emission with an integration time of 0.2 seconds per frame under the same above-band laser excitation scheme. We observe some wobbling of the emission line, which could attributed to the unstable charge environment and noise due to photo-excited free carriers under the above-band excitation. For a qualitative analysis of this fluctuation, we perform a fitting on the corresponding spectra, which reveals a fluctuation of 0.5 meV and an average central position of 1.5365 eV. This limitation needs to be addressed, as it could act as a bottleneck for indistinguishability as one of the potential sources of unwanted inhomogeneous broadening. To overcome the problem, several strategies could be adopted, for instance, the use of an encapsulation agent (e.g., hBN Parto et al. (2021)), electrostatic gating Lenferink et al. (2022), or employing advanced excitation schemes, i.e., resonant and SUPER scheme Karli et al. (2022). Moreover, in our out-of-plane electric contacts could also be employed to stabilize the charge environment and simultaneously provide a degree of tunability of the source via the Stark effect Noh et al. (2018).
To assess the purity of the exemplary SPE, we performed the second-order correlation measurement under the aforementioned 80 MHz pulsed excitation scheme, using a 750 nm LPF to isolate the main emission line and a typical HBT setup (Figure 5d). We observe an overlap of the peaks at non-zero delay times, separated by the laser repetition time of 12.5 ns, resulting from the increased lifetime of the emitter. More importantly, the significantly diminished coincidences at zero time delay indicate the high purity of single photons emitted by the SPE. To quantify this figure of merit, we employ a double-sided exponential function fitting for the adjacent peaks, which yields a value of . This result demonstrates the exceptional single-photon purity that can be achieved through the proposed fabrication novelties on WSe monolayer quantum emitters.
Given the showcased control over the degree of polarization of single photons and spatial localization of the emitters, future works could integrate the SPEs into nanophotonic cavities to enhance the radiative emission rate via the Purcell effect. Decreasing the spontaneous emission lifetime of our nanowrinkle emitters would contribute to both increasing the photon indistinguishability Liu et al. (2018) and efficiently coupling the emitted photons to the optical cavity mode Gregersen et al. (2013). Moreover, in-depth studies on the injection of electric and magnetic fields could contribute to a more comprehensive understanding of the fundamental mechanisms governing single-photon emission in these materials and provide us with spectrally stable sources by minimizing charge noise. This could lead to the introduction of tunability over key characteristics of the emitters (e.g., main emission wavelength). Theoretical investigations may focus on exciton-phonon interactions, providing crucial insights into the advancement of SPEs in TMD platforms. At the same time, advanced excitation schemes like the SUPER scheme Bracht et al. (2021); Karli et al. (2022) could potentially minimize the interaction with phonons, resulting in a narrower ZPL.
III Conclusion
In conclusion, we presented an efficient method for generating highly polarized single photons from orientation-controlled nanowrinkles in monolayer WSe. This method employs novel nanostructures to deterministically induce one-directional strain, enabling precise control of the polarization of emitted single photons. We have successfully demonstrated single photon emission from WSe nanowrinkle emitters with high purity () and a high degree of polarization (99 4%). These findings provide crucial insights for the future integration of these quantum emitters into a specific mode of desired photonic structures, e.g., the TE mode of a planar ridge waveguide. Understanding the fabrication limitations and critical aspects of deterministic polarization control of quantum emitters are essential steps toward their effective utilization as functional single-photon sources in quantum photonics.
IV Methods
IV.1 Sample preparation
Nanopillar fabrication: The substrate was prepared by thermally oxidizing 110 nm of SiO on a silicon wafer. Alignment marks were patterned by UV lithography (using the AZ 5214E positive resist) and subsequent electron-beam evaporation of 5 nm of Ti and 50 nm of Au. After cleaving the wafer into chips, the fabrication of the nanopillars was carried out by spin coating a high-resolution negative e-beam resist (hydrogen silsesquioxane or HSQ, XR-1541-006) at 3000 rpm for 1 min, followed by two soft baking steps at 120 C and 220 C, both for 2 minutes. The resist was patterned by an e-beam writer (JEOL 9500, 100 kV, 6 nA) with a dose of 11000 µC/cm and developed in a 1:3 solution of AZ 400K : HO to obtain 150 nm-tall glass nanopillars in four different shapes, namely cylindrical pillars, three-pointed stars, five-pointed stars, and bowties. It is worth remarking that a high dose is required to cure the spin-on dielectric, whose composition becomes SiO. Other negative e-beam resists with high sensitivity (activated at doses as low as 30 µC/cm) have also been tested (Supporting Note S5).
Exfoliation and transferring: WSe2 flakes were first exfoliated from bulk crystals (HQ Graphene) by using the scotch-tape method Castellanos-Gomez et al. (2014). The flakes were then exfoliated on a polydimethylsiloxane (PDMS) stamp, which was prepared on a glass slide. The monolayers were identified through their PL emission at RT by using a 450 nm LED source and were then dry-transferred on the nanostructures with a transfer stage, which allows heating the chip to 70 C. At this temperature, the van der Waals interactions between the flake and the substrate overcome the adhesion to PDMS, and the flake is released.
Defect fabrication: Defects were introduced on the monolayer lattice by bombarding the strained areas of the material with an electron beam (JEOL 9500 e-beam writer, 100 kV, 1000 µC/cm).
IV.2 Optical characterization
Photoluminescence measurements: The PL measurements were obtained with a custom-built low-temperature micro-photoluminescence (µPL) setup. The sample is mounted on a closed-cycle cryostat (attoDRY800) operating at a base temperature of 4 K. The cryostat is equipped with piezoelectric nanopositioners and a low-temperature microscope objective ( NA = 0.82) located inside the cryostat. The sample was excited with an LED at 470 nm for the continuous-wave (CW) excitation and an 80 MHz femtosecond laser (Chameleon Ultra II, Coherent) at 550 nm for the pulsed excitation. The PL spectra were acquired with a fiber-coupled spectrometer (iHR 550, Horiba). A polarizer mounted on a motorized rotation mount was placed on the collection path for polarization-resolved measurements. PL images were acquired with an integration time of 1 s using a CMOS camera (Tucsen Dhyana, 4 MP).
Second-order correlation measurements: For second-order correlation measurements, we employed a fiber-optic Hanbury-Brown and Twiss (HBT) interferometer. The photon counting was conducted using superconducting nanowires single-photon detectors (ID218, ID Quantique) connected to a time controller (ID900, IDQ).
IV.3 Sample imaging
Bright-field (BF) images of the sample were taken using a confocal-microscopy setup equipped with an objective lens (Nikon LU Plan Fluor, , NA = 0.8) and a CMOS camera (CS165MU/M, Thorlabs).
The topographic images of the samples were obtained with a scanning electron microscope (SEM) (Zeiss Supra 40 VP, 5 kV, SE2 detector). To avoid creating further defects in the flakes, SEM imaging was carried out after all the measurements for optical characterization.
Finally, the surface morphology of the samples, namely the width and height of the nanopillars and the WSe nanowrinkles, was characterized by atomic force microscopy (Bruker Dimension Icon-PT, ScanAsyst-air tip).
Associated Content
Supporting Information
The Supporting Information is available free of charge at https://xxxxxxxx.xxxxx.
Figure S1: Schematic representation of the fabrication flow, Figure S2: Comparison of PL spectra before and after e-beam irradiation, Figure S3: Investigation on the morphological profile of the nanowrinkles, Figure S4: SEM images of TS and FS structures with a statistical analysis on the misalignment of the nanowrinkles, Figure S5: SEM images of BT structures with a histogram of misalignment of the nanowrinkles, Figure S6: Comparison of the PL emission of an exemplary quantum emitter under two excitation schemes (LED, 470 nm and pulsed laser, 532 nm), Figure S7: AFM image of a WSe monolayer flake on a TS structure made from AR-N 7520.11.
Funding
The authors acknowledge support from the European Research Council (ERC-StG “TuneTMD”, grant no. 101076437) and the Villum Foundation (grant no. VIL53033). The authors also acknowledge the European Research Council (ERC-CoG“Unity”, grant no.865230). NG acknowledges support from the European Union’s Horizon 2020 Research and Innovation Program under the Marie Skolodowska-Curie Grant Agreement no. 861097.
Notes
The authors declare no competing financial interest.
Acknowledgement
The authors acknowledge the cleanroom facilities at DTU Nanolab – National Centre for Nano Fabrication and Characterization.
Author contributions
AP fabricated the quantum emitter samples. CP and AM supported the project by fabricating various shaped pillars for strain engineering. AP, CP, AM, and BM performed optical measurements from the fabricated samples. AP, PM, and CP contributed to data analysis and processing. BM conceived the idea, and LV, NG, and BM coordinated the project. AP, PM, and BM wrote the manuscript with input from all co-authors.